# Market Beliefs about Open vs. Closed AI

### v1

@@ Results
For inflation protected bonds (TIPS), the movement on both open and closed releases are less extreme, shown in Figure 11 and Table 4: for 30 year, a estimated decrease of 7-10 bps following closed releases (p-values 0.03-0.16) and increase of 7-9 bps following open releases (p-values 0.08-0.19). There remains a difference but it is narrower, 15-18 bps (p-values 0.01-0.04). Corroborating this, Figure 5 panel (a) finds that inflation expectations move up on the release of open models, and down on the release of closed models. (Figure 12 shows that increases in inflation expectations coincide with increased gaps between treasuries and TIPS.)
@@
I plot the path of yields for each individual event in Figures 14, 15, and 16. The entire distribution of paths is shifted between open and closed releases for longer maturity bonds, and the pattern is not driven by outliers or the releases of particular labs.
@@ 3.1.1 Alternate dimensions of heterogeneity
Yet it would be reasonable to expect a dose response, whereby larger shifts in expectations correlate with larger changes in yields. However, if open and closed releases induce opposing effects, large positive expectation shifts could have counterveiling effects for the two types of releases. Figure 20 thus considers the interaction. The point estimates suggest that releases that induce forecasters to anticipate sooner arrival of AGI have a more positive effect on 30-year treasury yields for open releases, and a stronger negative effect for closed releases. However, these estimates are extremely imprecise.777They also do not account for overlap with the categories omitted from the plot.
These results suggest that these associated factors are less likely to explain the opposing shifts.
@@ Discussion
There is wide disagreement about the economic implications of advances in machine intelligence. This paper builds on a literature that attempts to better understand market beliefs about these implications through financial decisions (jones_ai_2024; chow_transformative_2025; andrews_markets_2025). This paper investigates the puzzle documented by andrews_markets_2025. It finds evidence that long term bond markets shift in opposite directions following open versus closed AI model releases. The patterns documented here could be explained by many different stories. One possibility is that firms make different investments depending on the two types of advancements. While established tech firms have substantial internal resources to fund AI investments, open models allow for wider categories of firms to invest in developing and integrating AI technologies. It is possible that this could increase demand for resources in the present. However, it could also be that markets believe that openness leads to higher economic growth or higher existential risk than closed models. Regardless of the underlying mechanism, these results suggest the possibility that market participants believe that the openness of AI models could have large economic implications.

### v3

@@ Results
Estimates of cumulative returns. Cumulative returns from 15 days prior to release, using equation (1). Constant maturity duration noted in years. 90% and 95% confidence intervals are shaded.
Results for all events pooled for US treasury bonds are shown in Figure 3, given the estimates of equation (1). As suggested by Figure 1, the yield impact is smaller and not statistically significant from zero when open and closed releases are pooled. The remainder of the paper focuses on the split between open and closed releases.
@@
For corporate bonds above 3 year duration, results are similar, as shown in Figure 11 and Table 3. For inflation protected bonds (TIPS), the movement on both open and closed releases are less extreme, shown in Figure 12 and Table 4: for 30 year, a estimated decrease of 7-10 bps following closed releases (p-values 0.03-0.16) and increase of 7-9 bps following open releases (p-values 0.08-0.19). There remains a difference but it is narrower, 15-18 bps (p-values 0.01-0.04). Corroborating this, Figure 6 panel (a) finds that inflation expectations move up on the release of open models, and down on the release of closed models. (Figure 13 shows that increases in inflation expectations coincide with increased gaps between treasuries and TIPS.)
@@
I plot the path of yields for each individual event in Figures 15, 16, and 17. The entire distribution of paths is shifted between open and closed releases for longer maturity bonds, and the pattern is not driven by particular releases.
@@
The divergence is also clear when events are left out using a jackknife test, though for some subsamples the difference loses statistical significance. Regression results are shown in Appendix Figure 21 (dropping one pair of events; one from open and one from closed) and Appendix Figure 22 (dropping two pairs of events). Median changes are shown in Appendix Figure 23 (one pair) and Figure 24 (two pairs).
The divergence is also clear under different window lengths. It is similar for day windows around the events, using OLS (Figure 25) or the median change estimator (Figure 26). Under these longer windows, the yield paths are approximately zero until approximately day -15, which supports the window length of the main specification. A smaller divergence is present in a day windows around the events, whether using OLS (Figure 27) or the median change estimator (Figure 28). In those specifications it is statistically harder to distinguish the yield path of open releases from trend, but the difference between the paths of open and closed releases and the path of closed releases are relatively less likely in the placebo distributions.
Another robustness exercise controls for news that may have happened simultaneously, using the residualization procedure of andrews_markets_2025, in Appendix A.2. AI releases may themselves contribute to these indices, but nonetheless I find that it can be harder to distinguish the yield path of open releases from trend but the magnitude remains large, and the difference between open and closed releases remains large and statistically significant.
@@ 3.1.1 Alternate dimensions of heterogeneity
It would be reasonable to expect that larger shifts in expectations correlate with larger changes in yields. However, if open and closed releases induce opposing effects, large positive expectation shifts could have counterveiling effects for the two types of releases. Figure 29 thus considers the interaction. The point estimates suggest that releases that induce forecasters to anticipate sooner arrival of AGI have a very slightly more positive effect on 30-year treasury yields for open releases, and a stronger negative effect for closed releases. However, these estimates are extremely imprecise.777They also do not account for overlap with the categories omitted from the plot.
These results suggest that these associated factors are less likely to explain the opposing shifts. However, it is possible that these performance measures imperfectly capture the economic importance of a model, and that the heterogeneity I observe is driven by some other associated factor.
@@ Discussion
There is wide disagreement about the economic implications of advances in machine intelligence. This paper builds on a literature that attempts to better understand market beliefs about these implications through financial decisions (jones_ai_2024; chow_transformative_2025). This paper investigates the puzzle documented by andrews_markets_2025. It finds evidence that long term bond markets shift in opposite directions following open versus closed AI model releases. The patterns documented here could be explained by many different stories. One possibility is that firms make different investments depending on the two types of advancements. While established tech firms have substantial internal resources to fund AI investments, open models allow for wider categories of firms to invest in developing and integrating AI technologies. It is possible that this could increase demand for resources in the present. It could also result from other economic decisions impacted by whether advances in AI are open or closed. It could also be that markets believe that open advances lead to higher economic growth or higher existential risk than closed advances. Regardless of the underlying mechanism, these results suggest the possibility that market participants believe that whether advances in AI are open or closed could have important economic implications.
@@ Appendix A Appendix
Controlling for other news
As an additional robustness check, I follow andrews_markets_2025 in residualizing the main series based on other news, with the Citigroup US Economic Surprise Index (citigroup_global_markets_citigroup_2025), the VIX volatility index, and the Federal Reserve Bank of San Francisco Daily News Sentiment Index (shapiro_measuring_2022).
I first regress the yield outcome series on a constant plus 15 lags of each control series, and then run the regression procedure detailed in Section 3. I conduct this for the level of surprise (Figure 33), daily changes in VIX (Figure 34), level of news sentiment (Figure 35), and all three (Figure 36). In some of these, the confidence bands for open include zero, but the magnitudes remain large and positive, and the difference between open and closed releases remains large and statistically significant.
@@
OLS coefficient estimates. The shaded regions represent 90% and 95% of a placebo distribution, over 5,000 permutation draws. Dashed line indicates mean of the placebo distribution. Dotted line represents the main estimate. Thin lines represent the estimate resulting from dropping each pair of events (one from the open sample and one from the closed sample). For the first column, the placebo includes all dates in data. For last column, the placebo distribution includes the pooled event dates of closed and open AI model releases.
OLS coefficient estimates. The shaded regions represent 90% and 95% of a placebo distribution, over 5,000 permutation draws. Dashed line indicates mean of the placebo distribution. Dotted line represents the main estimate. Thin lines represent the estimate resulting from dropping two pairs of events (two from the open sample and two from the closed sample). For the first column, the placebo includes all dates in data. For last column, the placebo distribution includes the pooled event dates of closed and open AI model releases.
Dotted line shows the estimated change for the median event, relative to 15 days before the event. The shaded regions represent 90% and 95% of a placebo distribution, over 5,000 permutation draws. Dashed line indicates mean of the placebo distribution. Thin lines represent the estimate resulting from dropping each pair of events (one from the open sample and one from the closed sample). For the first column, the placebo includes all dates in data. For last column, the placebo distribution includes the pooled event dates of closed and open AI model releases.
Dotted line shows the estimated change for the median event, relative to 15 days before the event. The shaded regions represent 90% and 95% of a placebo distribution, over 5,000 permutation draws. Dashed line indicates mean of the placebo distribution. Thin lines represent the estimate resulting from dropping two pairs of events (two from the open sample and two from the closed sample). For the first column, the placebo includes all dates in data. For last column, the placebo distribution includes the pooled event dates of closed and open AI model releases.
Solid line shows the estimated change for the median event, relative to 30 days before the event. The shaded regions represent 90% and 95% of a placebo distribution, over 5,000 permutation draws. Dashed line indicates mean of the placebo distribution. For left column, the placebo includes all dates in data. For right column, the placebo distribution includes the pooled event dates of closed and open AI model releases.
Solid line shows the estimated change for the median event, relative to 5 days before the event. The shaded regions represent 90% and 95% of a placebo distribution, over 5,000 permutation draws. Dashed line indicates mean of the placebo distribution. For left column, the placebo includes all dates in data. For right column, the placebo distribution includes the pooled event dates of closed and open AI model releases.
@@
Estimates of cumulative returns, residualized against the Citigroup US Economic Surprise Index, as described in Appendix A.2. The left panels show cumulative returns from 15 days prior to release, using equation (2). The right panels show the estimated difference. Constant maturity duration noted in years. 90% and 95% confidence intervals are shaded.
Estimates of cumulative returns, residualized against VIX, as described in Appendix A.2. The left panels show cumulative returns from 15 days prior to release, using equation (2). The right panels show the estimated difference. Constant maturity duration noted in years. 90% and 95% confidence intervals are shaded.
Estimates of cumulative returns, residualized against the Federal Reserve Bank of San Francisco Daily News Sentiment Index, as described in Appendix A.2. The left panels show cumulative returns from 15 days prior to release, using equation (2). The right panels show the estimated difference. Constant maturity duration noted in years. 90% and 95% confidence intervals are shaded.
Estimates of cumulative returns, residualized against the Citigroup US Economic Surprise Index, VIX, and the Federal Reserve Bank of San Francisco Daily News Sentiment Index, as described in Appendix A.2. The left panels show cumulative returns from 15 days prior to release, using equation (2). The right panels show the estimated difference. Constant maturity duration noted in years. 90% and 95% confidence intervals are shaded.